# Concurrent Composition for Differentially Private Continual Mechanisms

### v2

@@ A.3. Construction of $\phi_{t}$
The proof consists of two steps: (i) ensuring that the optimization problems are feasible, and (ii) applying the transfinite recursion theorem (Lemma A.26) and proving that $\phi_{t}$ exists.
Fix any $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$. By Remark A.24, each answer in $A$ is either the smallest element of $A$ or has an immediate predecessor. We will inductively show that for every $a_{t}\in A$, the values $\frac{1}{\delta}\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})$ and $\frac{1}{\delta}\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})$ satisfy the constraints of the optimization problem defining $\phi_{t}^{0}(q_{1},\dots,a_{t})$ and $\phi_{t}^{1}(q_{1},\dots,a_{t})$.
$\frac{1}{\delta}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t})$ trivially satisfies constraint ($\phi$1). Moreover, by Corollary A.17, it satisfies constraint ($\phi$3). To prove constraints ($\phi$2) is satisfied, we need to show that
(2) $\sum_{a\in\Gamma_{<a_{t}}}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ $+\sum_{a\in\Gamma_{>a_{t}}\cup\{a_{t}\}}\frac{1}{\delta}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}),$
We prove this for the base case (i.e., when $a_{t}$ is the smallest element in $A$) and the inductive step (i.e., when $a_{t}$ has a predecessor $a^{*}$) separately:
@@
We know that $\phi_{t-1}(q_{1},\dots,a_{t-1})$ satisfies condition (I). Thus, by Lemma A.21,
finishing the proof for the base case.
Inductive Step: Let $a_{*}$ be the immediate predecessor of $a_{t}$. Since $\Gamma_{<a_{t}}=\Gamma_{<a^{*}}\cup\{a^{*}\}$, constraints ($\phi$2) for defining $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ is identical to (2). By inductive hypothesis, $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ exists and satisfies these constraints. Thus, inequality (2) holds, completing the induction.
Since $<_{t}$ is a well-order and the value of $\phi_{t}(q_{1},\dots,a_{t})$ depends only on preceding values in this order, by the transfinite recursion theorem, the function $\phi_{t}$ is well-defined and exists. ∎
A.3.3. Satisfying Conditions (I) to (III)
Our main objective is to construct $\phi_{t}^{b}$ so that it fulfills conditions (I), (II), and (III). Constraints ($\phi$1) and ($\phi$3) directly imply conditions (I) and (II), respectively. It remains to show $\phi_{t}$ satisfies conditions (III).
First, we claim for each $b\in\{0,1\}$ and every transcript $(q_{1},\dots,a_{t-1},q_{t})$, $\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ is at most $\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$: By constraint ($\phi$2) and non-negativity of $\mathrm{L}^{b}$, for every $a_{t}$,
$\sum_{a\in\Gamma_{<a_{t}}\cup\{a_{t}\}}\phi_{t}^{b}(q_{1},\dots,a_{t})\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
Thus, when $A$ is finite and $a^{\prime}$ is the maximum element of $A$, we have
$\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)=\sum_{\mathclap{a\in\Gamma_{<a^{\prime}}\cup\{a^{\prime}\}}}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a^{\prime})\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
Moreover, when $A$ is infinite and $a_{1}^{\prime}<_{A}a_{2}^{\prime}<_{A}\dots$ are the sorted elements of $A$, we have
$\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ $=\lim_{k\to\mathbb{N}}\sum_{a\in\Gamma_{<a_{k}^{\prime}}\cup\{a_{k}^{\prime}\}}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a_{k}^{\prime})\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
The rest of proof is identical to the proof given in [18]. We repeat it for completeness. For the sake of contradiction, suppose there exists $b\in\{0,1\}$ and $(q_{1},\dots,a_{t-1},q_{t})$ such that
@@ A.3.3. Satisfying Conditions (I) to (III)
Fix $b$. As $\phi_{t}^{b}$ is lower bounded by $\mathrm{L}_{t}^{b}/\delta$, this implies that for every $a_{t}\in A$, in the optimization problem determining the value of $\phi_{t}(q_{1},\dots,a_{t})$, constraint ($\phi$2) for $b$ is not tight. Consequently, for every $a_{t}\in A$, we must have had
$\phi_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\phi_{t}^{1-b}(q_{1},\dots,a_{t})=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$
otherwise, the value of $\phi_{t}^{b}(q_{1},\dots,a_{t})$ could have increased, which contradicts the maximality of $\phi^{t}_{b}$.
Taking a sum over all $a_{t}\in A$, we get
@@
By (II) and the inductive hypothesis, we know that the RHS is at least
@@
$\sum_{a_{t}\in A}\phi_{t}^{b}(q_{1},\dots,a_{t})$ $-\phi_{t}^{b}(q_{1},\dots,a_{t-1})\geq e^{-\epsilon}(\sum_{a_{t}\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t})-\phi_{t}^{1-b}(q_{1},\dots,a_{t-1}))$
@@
$\sum_{a_{t}\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t})$ $-\phi_{t}^{1-b}(q_{1},\dots,a_{t-1})\geq e^{-\epsilon}(\sum_{a_{t}\in A}\phi_{t}^{b}(q_{1},\dots,a_{t})-\phi_{t}^{b}(q_{1},\dots,a_{t-1})),$

### v3

@@ 8.3. Construction of $\phi_{t}$
The proof consists of two steps: (i) proving that $\mathcal{C}_{q_{1},\dots,a_{t}}$ is non-empty, and (ii) applying the transfinite recursion theorem (Lemma 8.25) and concluding that $\phi_{t}$ exists.
(i): Fix any $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$. By Remark 8.23, each answer in $A$ is either the smallest element of $A$ or has an immediate predecessor. We will inductively show that for every $a_{t}\in A$, the values $\frac{1}{\delta}\xi_{t}^{0}(q_{1},\dots,a_{t})$ and $\frac{1}{\delta}\xi_{t}^{1}(q_{1},\dots,a_{t})$ satisfy all the constraints for the optimization problem defining $\phi_{t}(q_{1},\dots,a_{t})$, and consequently, $\mathcal{C}_{q_{1},\dots,a_{t}}$ is not empty.
First, note that $\frac{1}{\delta}\xi_{t}^{b}(q_{1},\dots,a_{t})$ trivially satisfies Constraint ($\phi$.1). Moreover, by Constraint ($\xi$.3), it satisfies Constraint ($\phi$.3). To prove Constraint ($\phi$.2) is satisfied,it suffices to show that
(3) $\sum_{a\in\Gamma_{<a_{t}}}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ $+\sum_{a\in\Gamma_{>a_{t}}\cup\{a_{t}\}}\frac{1}{\delta}\xi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}),$
We prove this by induction on $a_{t}\in A$:
@@
$\sum_{\begin{subarray}{c}a\in A\end{subarray}}\xi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
Let $a+t=a_{1}^{\prime}<_{A}a_{2}^{\prime}<_{A}\dots$ denote the ordered elements of $A$. As $\mathrm{L}_{t}$ is non-negative, by Constraint ($\xi$.2), for every $n\in\mathbb{N}$, we have
$\sum_{\begin{subarray}{c}a\in\Gamma_{<a_{n}^{\prime}}\cup\{a_{n}^{\prime}\}\end{subarray}}\xi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
Taking $\lim_{n\to\infty}$ when $A$ is infinite or setting $n=|A|$ when $A$ is finite gives inequality (3).
Inductive Step: Let $a^{*}$ be the immediate predecessor of $a_{t}$. Since $\Gamma_{<a_{t}}=\Gamma_{<a^{*}}\cup\{a^{*}\}$, Constraints ($\phi$.2) for defining $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ is identical to (3). By inductive hypothesis, $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ exists and satisfies ($\phi$.2). Thus, inequality (3) holds, completing the induction.
(ii): For every $(q_{1},\dots,a_{t-1},q_{t})$, the values of the function $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ is uniquely determined by the preceding values of this function in the well-order $_{A}$. Thus, by the transfinite recursion theorem, the function $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ exists and is uniquely defined. Consequently, the function $\phi_{t}$ exists and is uniquely defined. ∎
8.3.4. Satisfying Conditions (I) to (IV)
Our objective was to construct $\phi_{t}^{b}$ so that it fulfills conditions (I) to (IV). Combination of Constraints ($\phi$.1) and ($\xi$.1) gives condition (I). Moreover, Constraint ($\phi$.3) directly imply condition (II). The proof of the following claim is identical to a proof in [19]:
Claim 8.31.
$\phi_{t}$ satisfies conditions (III).
Proof.
For each $b\in\{0,1\}$ and every transcript $(q_{1},\dots,a_{t-1},q_{t})$, by Constraint ($\phi$.2) and non-negativity of $\xi^{b}$, for every $a_{t}$,
$\sum_{a\in\Gamma_{<a_{t}}\cup\{a_{t}\}}\phi_{t}^{b}(q_{1},\dots,a_{t})\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}),$
which in both cases where $A$ is finite and countable implies that
$\sum_{a\in A}\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}).$
For the sake of contradiction, suppose there exists $b\in\{0,1\}$ and $(q_{1},\dots,a_{t-1},q_{t})$ such that
@@ 8.3.4. Satisfying Conditions (I) to (IV)
Fix $b$. As $\phi_{t}^{b}$ is lower bounded by $\xi_{t}^{b}/\delta$ and the values of $\phi_{t}$ are maximal solutions, this strict inequality implies that for every $a_{t}\in A$, Constraint ($\phi$.2) for $b$ is not tight. Consequently, for every $a_{t}\in A$, we must have had
$\delta\phi_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\delta\phi_{t}^{1-b}(q_{1},\dots,a_{t})=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$
otherwise, the value of $\phi_{t}^{b}(q_{1},\dots,a_{t})$ could have increased, which contradicts the maximality. Taking a sum over all $a_{t}\in A$, we get
@@
By (II) and Remark 8.21, we know that the RHS is at least
@@
$\sum_{a_{t}\in A}\phi_{t}^{b}(q_{1},\dots,a_{t})$ $-\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})\geq e^{-\epsilon}(\sum_{a_{t}\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t})-\phi_{t-1}^{1-b}(q_{1},\dots,a_{t-1}))$
@@
$\sum_{a_{t}\in A}\phi_{t}^{1-b}(q_{1},\dots,a_{t})$ $-\phi_{t-1}^{1-b}(q_{1},\dots,a_{t-1})\geq e^{-\epsilon}(\sum_{a_{t}\in A}\phi_{t}^{b}(q_{1},\dots,a_{t})-\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})),$